In a town, the probabilities that a family owns a microwave, an air conditioner and a computer are 0.5, 0.6, and 0.8 respectively.

A)If a family is chosen at random from the town, find the probability that the family chosen owns

(i) all three items

(ii) only a microwave and an air conditioner

(iii) none of the three it


1
Expert's answer
2021-03-31T07:16:01-0400

i) p = 0.5 "\times" 0.6 "\times" 0.8 = 0.24

ii) p = 0.5 "\times" 0.6 "\times" (1 - 0.8) = 0.06

iii) p = (1 - 0.5) "\times" (1 - 0.6) "\times" (1 - 0.8) = 0.04